# SAFER-Splat: A Control Barrier Function for Safe Navigation with Online Gaussian Splatting Maps

### v1

@@ Abstract
SAFER-Splat (Simultaneous Action Filtering and Environment Reconstruction) is a real-time, scalable, and minimally invasive action filter, based on control barrier functions, for safe robotic navigation in a detailed map constructed at runtime using Gaussian Splatting (GSplat). We propose a novel Control Barrier Function (CBF) that not only induces safety with respect to all Gaussian primitives in the scene, but when synthesized into a controller, is capable of processing hundreds of thousands of Gaussians while maintaining a minimal memory footprint and operating at 15 Hz during online Splat training. Of the total compute time, a small fraction of it consumes GPU resources, enabling uninterrupted training. The safety layer is minimally invasive, correcting robot actions only when they are unsafe. To showcase the safety filter, we also introduce SplatBridge, an open-source software package built with ROS for real-time GSplat mapping for robots. We demonstrate the safety and robustness of our pipeline first in simulation, where our method is 20-50x faster, safer, and less conservative than competing methods based on neural radiance fields. Further, we demonstrate simultaneous GSplat mapping and safety filtering on a drone hardware platform using only on-board perception. We verify that under teleoperation a human pilot cannot invoke a collision. Our videos and codebase can be found at https://chengine.github.io/safer-splat.
@@ I Introduction
Although algorithms for safe robotic control have been developed for a broad range of map representations, e.g., [15, 16, 17, 18, 14], these existing methods either require a pre-built map, making them unsuitable for online scenarios, or require stringent assumptions on the robot’s dynamics, sensing modalities, or its nominal controller. SAFER-Splat leverages a novel CBF, closely integrated with the GSplat representation, to derive a lightweight, minimally-invasive safe action filter. To synthesize a controller, we embed the CBF safety constraint into a quadratic program, minimizing the deviation between the desired and actuated control effort. By pruning the CBF constraints to identify a minimal number of constraints, our method scales efficiently to run in real-time with hundred of thousands ellipsoidal primitives in the scene. Together, the safety guarantees, computational efficiency, and scalability of SAFER-Splat enable its superior performance in collision-free robot control in online GSplat scenes. To showcase SAFER-Splat in a full autonomy stack, we also introduce SplatBridge, a ROS bridge for real-time, online GSplat training for robots.
@@ II Related Work
In this section, we review the related literature in radiance fields, SLAM using radiance fields, robot control using barrier functions, and planning and control in radiance fields.
Radiance Fields and SLAM. First introduced in [5], Neural Radiance Fields demonstrated the photorealistic novel view rendering capability of radiance field scene representations, where the environment is represented by spatial density and color fields, parameterized by multi-layer perceptrons (MLPs). In addition to detailed scene color modeling, NeRFs produce high-fidelity, dense geometric scene reconstructions making them a potential alternative map representation for SLAM pipelines. Methods like iMAP [21] and NICE-SLAM [22] simultaneously optimize the NeRF network weights and the robot/camera poses. Meanwhile, NeRF-SLAM [23] and NeRFBridge [24] combine existing visual odometry with online NeRF training for real-time 3D scene reconstruction.
GSplat [7], a successor to NeRFs, retains a similar philosophy on the parametric modeling of scene radiance fields. However, rather than using an MLP-based parametric model, GSplat represents the scene using ellipsoidal primitives with a tile-based rasterization technique for differentiable rendering. Compared to NeRFs, GSplat provides superior training speed and reconstruction accuracy. GSplat has also been integrated into SLAM pipelines [20, 25, 26, 27]. Specifically, we integrate our safety layer with SplatBridge, a GSplat extension of NerfBridge. Both are open-source, ROS integrated software packages for online training of radiance fields from robot sensors.
@@
Planning and Control in Radiance Fields. Using NeRFs as an underlying scene representation, NeRF-Nav [17] plans trajectories for differentially flat robots, minimizing a collision cost through gradient descent, but has no safety guarantees and is slow to converge. CATNIPS [18] converts the NeRF into a probabilistic voxel grid to plan safe paths parameterized as Bézier curves. Most structurally similar to our work, [19] (which we call NeRF-CBF) uses NICE-SLAM [22] to train a NeRF and uses the rendered depth map at sampled poses to enforce step-wise safety using a CBF. NeRF-CBF requires a pre-trained NICE-SLAM map, due to the slow execution time of online NICE-SLAM. Furthermore, the discrete CBF has weak safety guarantees, and combined with slow NeRF rendering, introduces scalability issues. SAFER-Splat addresses these limitations by performing online mapping using SplatBridge and is amenable to continuous-time dynamics models.
To our knowledge, Splat-Nav [14] is the only work to propose a planning algorithm for a robot using a GSplat map, although we note that prior work on mapping with Gaussian Mixture Models exists [36, 37]. Splat-Nav converts the GSplat representation into a union of ellipsoids and performs rigorous collision checking to generate smooth, safe trajectories. Similarly to Splat-Nav, we leverage the ellipsoidal interpretation of GSplats to demonstrate safety of our safety filter however, our safety filter is meant to intervene at the low level feedback control layer, whereas Splat-Nav provides a high-level trajectory planning layer.
@@
We now formally present our definition of the scene representation which will be used by the safety layer. A scene consists of a union of $K$ ellipsoids comprising the GSplat $\mathcal{G}=\{\mathcal{E}_{i}\}_{i=1}^{K}$, where each ellipsoid $\mathcal{E}_{i}$ satisfies the standard form: ${\mathcal{E}_{i}=\{x\in\mathbb{R}^{3}|(x-\mu_{i})^{T}\Sigma_{i}^{-1}(x-\mu_{i})\leq 1\}}$ for means $\mu_{i}$ and covariance $\Sigma_{i}$. We decompose the covariances based on the rotation and scaling parameters from the GSplat formulation: $\Sigma_{i}=R\bar{S}\bar{S}^{T}R^{T},$ where $\bar{S}=\sqrt{\chi_{3}^{2}(\gamma)}S$ represents the scales of Gaussian primitive $i$ multiplied by the $\gamma\%$th percentile of the chi-squared distribution. For the remainder of the text, we will only refer to the $\gamma\%$ confidence ellipsoids (where $\gamma$ corresponds to $1\Sigma$ in this work) and denote $S$ as their semi-major axes.
@@ IV Control Barrier Functions
CBFs frame safety guarantees in the lens of forward invariance. Given a CBF candidate ${h(x):\mathbb{R}^{n}\to\mathbb{R}}$ of the robot’s state $x\in\mathcal{D}$, we define the safe set as the 00{0}-superlevel set $\Omega=\{x\in\mathcal{D}\;|\;h(x)\geq 0\}$. The $\mathrm{int}(\Omega)$ is defined where $h(x)$ is strictly positive, whereas the boundary $\partial\Omega$ is precisely where $h(x)$ is 0. In order for a safety algorithm to be safe, we must remain in $\Omega$ for all time $t\in\mathbb{R}_{+}$. The forward invariance can be formalized through the condition [39] (an extension of Nagumo’s Theorem [40]): ${\dot{h}(x(t))\geq-\alpha(h(x))}$, where $\alpha$ is any class-$\kappa$ extended function. Intuitively, this constraint implies that the CBF candidate is allowed to decrease at an arbitrarily high rate until it reaches $\partial S$, at which point it can no longer be negative and become unsafe. Conversely, the constraint is also defined for unsafe behavior, in which the CBF is designed to increase $h(x)$ back into $\Omega$.
In order to synthesize a controller, we first define the continuous-time control-affine dynamics: ${\dot{x}=f(x)+g(x)u,}$ where state transition function ${f(x):\mathcal{D}\to\mathbb{R}^{n}}$, actuation matrix ${g(x):\mathcal{D}\to\mathbb{R}^{n\times m}}$, and control effort $u\in\mathbb{R}^{m}$. Due to space limitations, we limit our discussion to double-integrator systems and refer readers to [29] for a more detailed discussion of CBFs. We can define a formal CBF constraint for double-integrator systems: ${\forall x\in\mathcal{D},\;\exists\;u:}$
$\begin{split}&\left[\left(\mathcal{L}_{f}\mathcal{L}_{g}h\right)^{T}u\geq-\mathcal{L}_{f}^{2}h-\alpha(\mathcal{L}_{f}h)-\beta(\mathcal{L}_{f}h+\alpha(h))\right],\end{split}$ (1)
@@ V Collision Avoidance in GSplat Maps
For a feasible integration with GSplat mapping methods, a safety layer must be scalable to the number of Gaussians in a GSplat and be safe at all timesteps.
@@
Considering the case where the robot is in the safe set, from the first-order optimality conditions, we have:
$\begin{split}\frac{\partial}{\partial\lambda}L(\hat{y}^{\star},\lambda^{\star};\hat{\mathbf{p}})&=\hat{y}^{\star T}S^{-2}\hat{y}^{\star}-1=0\\ \frac{\partial}{\partial\hat{y}}L(\hat{y}^{\star},\lambda^{\star};\hat{\mathbf{p}})&=2(\lambda^{\star}S^{-2}+\mathcal{I})\hat{y}^{\star}-2\hat{\mathbf{p}}=0.\\ \end{split}$ (5)
@@
which represents a tractable CBF candidate for spherical robots with radius $r$ and desired buffer distance $\epsilon$ navigating within GSplat maps. Note that for $\epsilon=0$, $h_{i}(\mathbf{p})$ will be positive in free space, 0 when the robot body just intersects the ellipsoid, and negative when penetrating the ellipsoid. Solving 1 requires the derivatives of the CBF. By the Envelope theorem [42], we can treat the optimized variables as constants when taking the first derivative. Hence,
@@
$\begin{split}\mathbf{H}_{i}&=2\phi_{i}(\mathbf{I}-\frac{\partial\hat{y}^{*}(\hat{\mathbf{p}})}{\partial\hat{\mathbf{p}}})\\ &=2\phi_{i}\left[\mathbf{I}-\frac{\partial\hat{y}^{*}(\lambda,\hat{\mathbf{p}})}{\partial\hat{\mathbf{p}}}+\frac{\partial\hat{y}^{*}(\lambda,\hat{\mathbf{p}})}{\partial\lambda}\frac{\partial\lambda}{\partial\hat{\mathbf{p}}}\right]\\ &=2\phi_{i}\left[\mathbf{I}-\mathrm{\mathbf{diag}}(\frac{S_{ii}^{2}}{\lambda+S_{ii}^{2}})+\frac{\partial\lambda}{\partial q}\frac{\partial\hat{y}^{*}}{\partial\lambda}\frac{\partial q}{\partial\hat{\mathbf{p}}}\right]\end{split}$ (9)
@@
We can extend the sphere-to-ellipsoid CBF to an ellipsoid-to-ellipsoid CBF, which is applicable to non-spherical robots, by defining a diffeomorphism ${\psi:\mathbb{O}\to\mathbb{W}}$ that maps the original manifold $\mathbb{O}$ to an equivalent manifold $\mathbb{W}$ where the robot is spherical. We can utilize the sphere-to-ellipsoid CBF in $\mathbb{W}$ and map the control back to $\mathbb{O}$.
@@
First, assume that the robot geometry is an ellipsoid $\mathcal{E}_{r}$ in the original 3D space $\mathbb{O}$, with body-to-world rotation matrix $R_{r}$, scaling matrix $S_{r}$ and centroid $\mathbf{p}$, and the scene ellipsoid’s $\mathcal{E}_{i}$ is as previously defined. By defining the diffeomorphism ${\psi:\mathbb{O}\to\mathbb{W}}$, where ${\psi(x)=\bar{S}^{-1}R_{r}^{T}x}$, we transform the robot from an ellipsoid in $\mathbb{O}$ to a unit sphere in $\mathbb{W}$. We transform all the ellipsoids representing the environment from $\mathbb{O}$ to $\mathbb{W}$, with the covariance of ellipsoid $i$ given by: $\Sigma_{i}^{\mathbb{W}}=S_{r}^{-1}R_{r}R_{i}S_{i}S_{i}^{T}R_{i}^{T}R_{r}^{T}S_{r}^{-T}$. An eigendecomposition is performed on $\Sigma_{i}^{\mathbb{W}}$ to retrieve $R_{i}^{\mathbb{W}}$ and $S_{i}^{\mathbb{W}}$. Then, we can use our sphere-to-ellipsoid CBF Eq. 7 in $\mathbb{W}$, making sure our dynamics are also expressed in $\mathbb{W}$. Specifically, our double-integrator dynamics are unchanged between $\mathbb{O}$ and $\mathbb{W}$. We only need to ensure that the desired acceleration $\bar{u}^{\mathbb{W}}$ is expressed in $\mathbb{W}$, while the outgoing filtered $u^{\mathbb{O}}$ is expressed in $\mathbb{O}$. ∎
@@
We can write (1) as the following half-space constraints:
@@
Note that for all safe states $x$ such that $h_{i}(x)>0$, $\phi_{i}=1$ and $\frac{S_{ii}^{2}}{\lambda+S_{ii}^{2}}<1$. Moreover, $\frac{\partial\lambda}{\partial q}\frac{\partial y^{*}}{\partial\lambda}\frac{\partial q}{\partial\hat{\mathbf{p}}}$ is of the form $aBB^{T}$ where $a\in\mathbb{R}_{++}$, forming a PSD matrix. $\mathbf{H}_{i}$ is positive-definite, hence the RHS (11) is always negative and $u=-(\alpha+\beta)v$ is a feasible solution. ∎
Finding a closed-form feasible point further allows us to avoid the costly routine of finding interior points, which is required to prune polytope constraints. When the robot is in collision, the corresponding constraints (11) do not necessarily contain the origin. However, there are typically very few ellipsoids in collision at any one time. Therefore, we are free to prune constraints for ellipsoids that are not in collision and keep all in-collision constraints. Thus, collision and collision-free scenarios have approximately the same computation savings, which amounts to roughly pruning $99.9\%$ of Gaussians.

### v2

# A Control Barrier Function for Safe Navigation with Online Gaussian Splatting Maps

@@ Abstract
SAFER-Splat (Simultaneous Action Filtering and Environment Reconstruction) is a real-time, scalable, and minimally invasive safety filter, based on control barrier functions, for safe robotic navigation in a detailed map constructed at runtime using Gaussian Splatting (GSplat). We propose a novel Control Barrier Function (CBF) that not only induces safety with respect to all Gaussian primitives in the scene, but when synthesized into a controller, is capable of processing hundreds of thousands of Gaussians while maintaining a minimal memory footprint and operating at 15 Hz during online Splat training. Of the total compute time, a small fraction of it consumes GPU resources, enabling uninterrupted training. The safety layer is minimally invasive, correcting robot actions only when they are unsafe. To showcase the safety filter, we also introduce SplatBridge, an open-source software package built with ROS for real-time GSplat mapping for robots. We demonstrate the safety and robustness of our pipeline first in simulation, where our method is 20-50x faster, safer, and less conservative than competing methods based on neural radiance fields. Further, we demonstrate simultaneous GSplat mapping and safety filtering on a drone hardware platform using only on-board perception. We verify that under teleoperation a human pilot cannot invoke a collision. Our videos and codebase can be found at https://chengine.github.io/safer-splat.
@@ I Introduction
Although algorithms for safe robotic control have been developed for a broad range of map representations, e.g., [15, 16, 17, 18, 14], these existing methods either require a pre-built map, making them unsuitable for online scenarios, or require stringent assumptions on the robot’s dynamics, sensing modalities, or its nominal controller. SAFER-Splat leverages a novel CBF, closely integrated with the GSplat representation, to derive a lightweight, minimally-invasive safety filter. To synthesize a controller, we embed the CBF safety constraint into a quadratic program, minimizing the deviation between the desired and actuated control effort. By pruning the CBF constraints to identify a minimal number of constraints, our method scales efficiently to run in real-time with hundred of thousands ellipsoidal primitives in the scene. Together, the safety guarantees, computational efficiency, and scalability of SAFER-Splat enable its superior performance in collision-free robot control in online GSplat scenes. To showcase SAFER-Splat in a full autonomy stack, we also introduce SplatBridge, a ROS bridge for real-time, online GSplat training for robots.
@@
The contributions of SAFER-Splat can be summarized as: (1) a tractable CBF-based safety filter for safe robot navigation in GSplat maps and (2) SplatBridge, a real-time ROS bridge for training GSplats online on robot platforms.
@@ II Related Work
Radiance Fields and SLAM. First introduced in [5], Neural Radiance Fields demonstrated the photorealistic novel view rendering capability of radiance field scene representations, where the environment is represented by spatial density and color fields, parameterized by multi-layer perceptrons. In addition to detailed scene color modeling, NeRFs produce high-fidelity, dense geometric scene reconstructions making them a potential alternative map representation for SLAM pipelines. Prior work [21, 22] simultaneously optimize the NeRF network weights and the robot/camera poses. Meanwhile, NeRF-SLAM [23] and NerfBridge [24] combine existing visual odometry with online NeRF training for real-time 3D scene reconstruction.
GSplat [7], a successor to NeRFs, retains a similar philosophy on the parametric modeling of scene radiance fields. However, rather than using an MLP-based parametric model, GSplat represents the scene using ellipsoidal primitives with a tile-based rasterization technique for differentiable rendering. Compared to NeRFs, GSplat provides superior training speed and reconstruction accuracy. GSplat has also been integrated into SLAM pipelines [20, 25, 26, 27]. Specifically, we integrate our safety layer with SplatBridge, a GSplat extension of NerfBridge.
@@
Planning and Control in Radiance Fields. Using NeRFs as an underlying scene representation, NeRF-Nav [17] plans trajectories for differentially flat robots, minimizing a collision cost through gradient descent, but has no safety guarantees and is slow to converge. CATNIPS [18] converts the NeRF into a probabilistic voxel grid to plan safe paths parameterized as Bézier curves. Most structurally similar to our work, NeRF-CBF [19] uses NICE-SLAM [22] to train a NeRF and uses the rendered depth map at sampled poses to enforce step-wise safety using a CBF. It requires a pre-trained NICE-SLAM map, due to the slow execution time of online NICE-SLAM. Furthermore, the discrete CBF has weak safety guarantees, and combined with slow NeRF rendering, introduces scalability issues. SAFER-Splat addresses these limitations by performing online mapping using SplatBridge and is amenable to continuous-time dynamics models.
To our knowledge, Splat-Nav [14] is the only work to propose a planning algorithm for a robot using a GSplat map, although we note that prior work on mapping with Gaussian Mixture Models exists [36, 37]. Splat-Nav converts the GSplat representation into a union of ellipsoids and performs rigorous collision checking to generate smooth, safe trajectories. Similarly, we leverage the ellipsoidal interpretation of GSplats to demonstrate safety of our safety filter; however, our safety filter is meant to intervene at the low level feedback control layer, whereas Splat-Nav provides a high-level trajectory planning layer.
@@
We now formally present our definition of the scene representation which will be used by the safety layer. A scene consists of a union of $K$ ellipsoids comprising the GSplat $\mathcal{G}=\{\mathcal{E}_{i}\}_{i=1}^{K}$, where each ellipsoid $\mathcal{E}_{i}$ satisfies the standard form: ${\mathcal{E}_{i}=\{x\in\mathbb{R}^{3}|(x-\mu_{i})^{T}\Sigma_{i}^{-1}(x-\mu_{i})\leq 1\}}$ for means $\mu_{i}$ and covariance $\Sigma_{i}$. We decompose the covariances based on the rotation and scaling parameters from the GSplat formulation: $\Sigma_{i}=R\bar{S}\bar{S}^{T}R^{T},$ where $\bar{S}=\sqrt{\chi_{3}^{2}(\gamma)}S$ represents the scales of Gaussian primitive $i$ multiplied by the $\gamma\%$th percentile of the chi-squared distribution. For this work, we use the $1\sigma$ ellipsoid with semi-major axes denoted as $S$.
@@ IV Control Barrier Functions
We provide a brief introduction to control barrier functions (CBFs) for double-integrator systems and refer interested readers to [29] for a more detailed discussion of CBFs. CBFs frame safety guarantees in the lens of forward invariance. Given a CBF candidate ${h(x):\mathbb{R}^{n}\to\mathbb{R}}$ of the robot’s state $x\in\mathcal{D}$, we define the safe set as the 00{0}-superlevel set ${\Omega=\{x\in\mathcal{D}\;|\;h(x)\geq 0\}}$. The $\mathrm{int}(\Omega)$ is defined where $h(x)$ is strictly positive, whereas the boundary $\partial\Omega$ is precisely where $h(x)$ is 0. In order for a safety algorithm to be safe, we must remain in $\Omega$ for all time $t\in\mathbb{R}_{+}$. The forward invariance can be formalized through the condition [39] (an extension of Nagumo’s Theorem [40]): ${\dot{h}(x(t))\geq-\alpha(h(x))}$, where $\alpha$ is any class-$\kappa$ extended function. Intuitively, this constraint implies that the CBF candidate is allowed to decrease at an arbitrarily high rate until it reaches $\partial S$, at which point it can no longer be negative and become unsafe. Conversely, the constraint is also defined for unsafe behavior, in which the CBF is designed to increase $h(x)$ back into $\Omega$.
In order to synthesize a controller, we first define the continuous-time control-affine dynamics: ${\dot{x}=f(x)+g(x)u,}$ where state transition function ${f(x):\mathcal{D}\to\mathbb{R}^{n}}$, actuation matrix ${g(x):\mathcal{D}\to\mathbb{R}^{n\times m}}$, and control effort $u\in\mathbb{R}^{m}$.
We can define a formal CBF constraint for double-integrator systems: ${\forall x\in\mathcal{D},\;\exists\;u:}$
$\left[\left(\mathcal{L}_{f}\mathcal{L}_{g}h\right)^{T}u\geq-\mathcal{L}_{f}^{2}h-\alpha(\mathcal{L}_{f}h)-\beta(\mathcal{L}_{f}h+\alpha(h))\right],$ (1)
@@ V Collision Avoidance in GSplat Maps
Considering the case where the robot is in the safe set, the optimal variables (denoted with *) must satisfy first-order optimality conditions:
$\begin{split}\frac{\partial}{\partial\lambda}L(\hat{y}^{\star},\lambda^{\star};\hat{\mathbf{p}})&=\hat{y}^{\star T}S^{-2}\hat{y}^{\star}-1=0\\ \frac{\partial}{\partial\hat{y}}L(\hat{y}^{\star},\lambda^{\star};\hat{\mathbf{p}})&=2(\lambda^{\star}S^{-2}+\mathbf{I})\hat{y}^{\star}-2\hat{\mathbf{p}}=0.\\ \end{split}$ (5)
@@
which represents a tractable CBF candidate for spherical robots with radius $r$ and desired buffer distance $\epsilon$. $\gamma$ and $\epsilon$ encapsulate uncertainty in the 3D reconstruction process and in the position estimate of the robot, which should be selected based on domain-specific knowledge. Note that for $\epsilon=0$, $h_{i}(\mathbf{p})$ will be positive in free space, 0 when the robot body just intersects the ellipsoid, and negative when penetrating the ellipsoid.
Solving 1 requires the derivatives of the CBF. By the Envelope theorem [42], we can treat the optimized variables as constants when taking the first derivative. Hence,
@@
$\begin{split}\mathbf{H}_{i}&=2\phi_{i}\left[\mathbf{I}-\mathrm{\mathbf{diag}}(\frac{S_{ii}^{2}}{\lambda+S_{ii}^{2}})+\frac{\partial\lambda}{\partial q}\frac{\partial\hat{y}^{*}}{\partial\lambda}\frac{\partial q}{\partial\hat{\mathbf{p}}}\right]\end{split}$ (9)
@@
The sphere-to-ellipsoid CBF (7) can be extended to an ellipsoid-to-ellipsoid CBF, applicable to ellipsoidal robots.
@@
An ellipsoidal robot in space $\mathcal{W}$ can be converted to a sphere in space $\mathcal{O}$ through a linear transformation $A:\mathcal{W}\to\mathcal{O}$ consisting of a rotation and anisotropic scaling. The set of ellipsoids are closed under linear transformations. Therefore, in $\mathcal{O}$, the sphere-to-ellipsoid CBF (7) can be used. Any action synthesized using dynamics in $\mathcal{O}$ should be mapped back to dynamics in $\mathcal{W}$. ∎
@@
We can write (1) as half-space constraints for every Gaussian:
@@
Note that for all safe states $x$ such that $h_{i}(x)>0$, $\phi_{i}=1$ and $\frac{S_{ii}^{2}}{\lambda+S_{ii}^{2}}<1$. Moreover, $\frac{\partial\lambda}{\partial q}\frac{\partial y^{*}}{\partial\lambda}\frac{\partial q}{\partial\hat{\mathbf{p}}}$ is of the form $aBB^{T}$ where $a\in\mathbb{R}_{++}$, forming a PSD matrix. $\mathbf{H}_{i}$ is positive-definite, hence the RHS (11) is always negative and $u_{0}=-(\alpha+\beta)v$ is a feasible solution. ∎
Having access to a closed-form feasible point circumvents the costly routine of finding interior points, which is a required step when pruning redundant polytope constraints. However, when the robot is in collision, the corresponding constraints (11) are not necessarily satisfied with $u_{0}$. We observe that there are typically very few ellipsoids in collision at any one time. Therefore, although we maintain all in-collision constraints, we are free to prune constraints for ellipsoids that are not in collision. Thus, collision and collision-free scenarios have approximately the same computation savings, which amounts to roughly pruning $99.9\%$ of all Gaussians.
Corollary 1.
The quadratic program (2) can be extended to enforce velocity and acceleration limits.
Proof.
A velocity barrier ($h_{v}(x)=v_{max}^{2}-\lVert v\rVert^{2}\geq 0$) constraint evaluated at $u_{0}$ for any $\kappa\in\mathbb{R}_{++}$ is
$\underbrace{-2v^{T}u_{0}}_{\dot{h}(x,u_{0})}\geq-\kappa\cdot h_{v}(x)\leftrightarrow\left(\kappa-2(\alpha+\beta)\right)\lVert v\rVert^{2}\leq\kappa v_{max}^{2}.$ (12)
Note that the RHS of (12) is always satisfied when $(\kappa,\alpha,\beta)>\mathbf{0}$ and $h_{v}(x)\geq 0$, hence the velocity barrier is forward invariant. Consequently, given the same feasible solution and a desired limit on the acceleration $\lVert u\rVert^{2}\leq a_{max}^{2}$, the following inequality must be satisfied
$\lVert u_{0}\rVert^{2}=(\alpha+\beta)^{2}\lVert v\rVert^{2}\leq a_{max}^{2}\;\forall\;\lVert v\rVert\leq v_{max},$ (13)
hence $\alpha+\beta\leq\frac{a_{max}}{v_{max}}$. Therefore, the inclusion of velocity and acceleration limits into (2) does not compromise the feasibility or safety of the program since $u_{0}$ satisfies (11, 12, and 13) so long as $\alpha,\beta$ are chosen accordingly.